# Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation

### v2

@@ Abstract
Climate change communication on social media increasingly employs microtargeting strategies to effectively reach and influence specific demographic groups. This study presents a post-hoc analysis of microtargeting practices within climate campaigns by leveraging large language models (LLMs) to examine Facebook advertisements. Our analysis focuses on two key aspects: demographic targeting and fairness. We evaluate the ability of LLMs to accurately predict the intended demographic targets, such as gender and age group, achieving an overall accuracy of $88.55\%$. Furthermore, we instruct the LLMs to generate explanations for their classifications, providing transparent reasoning behind each decision. These explanations reveal the specific thematic elements used to engage different demographic segments, highlighting distinct strategies tailored to various audiences. Our findings show that young adults are primarily targeted through messages emphasizing activism and environmental consciousness, while women are engaged through themes related to caregiving roles and social advocacy. In addition to evaluating the effectiveness of LLMs in detecting microtargeted messaging, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. We assess disparities in accuracy and error rates across demographic groups using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality. Our findings indicate that while LLMs perform well overall, certain biases exist, particularly in the classification of senior citizens and male audiences. The analysis of thematic explanations uncovers recurring patterns in messaging strategies tailored to various demographic groups, while the fairness analysis underscores the need for more inclusive and unbiased targeting methods. By showcasing the efficacy of LLMs in dissecting and explaining targeted communication strategies and by highlighting fairness concerns, this study provides a valuable framework for future research aimed at enhancing transparency, accountability, and inclusivity in social media-driven climate campaigns.
@@ 2 Related Work
Microtargeting on social media platforms such as Facebook has been widely studied in the context of political campaigns Islam et al. (2023a); Capozzi et al. (2021); Serrano et al. (2020); Silva et al. (2020); Capozzi et al. (2020) and public health messaging Islam and Goldwasser (2022); Silva and Benevenuto (2021); Mejova and Kalimeri (2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups Islam et al. (2023b). However, the effectiveness and ethical implications of such targeted messaging remain under-explored, particularly in terms of how different demographic groups are engaged and whether biases are present in the targeting strategies.
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans Chowdhery et al. (2023); Kojima et al. (2022); Le Scao et al. (2022); Brown et al. (2020); Gilardi et al. (2023); De Paoli (2023); Dai et al. (2023); Chiang and Lee (2023); Ziems et al. (2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into Thematic Analysis (TA). Researchers have proposed various frameworks, including an LLM-in-the-loop approach Islam and Goldwasser (2024b, a); Dai et al. (2023), integrating GPT-3 with expert-designed codebooks Xiao et al. (2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes Gao et al. (2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer Wei et al. (2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs predictions which have been highlighted in recent literatures Anthis et al. (2024); Dai et al. (2024); Kumar et al. (2023); Li et al. (2023b); Wu et al. (2023); Bender et al. (2021), emphasizing the need for rigorous evaluation of biases and disparities in model performance across different groups. In this paper, we leverage LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Facebook. Besides, we develop a new set of themes and aspects based on those explanations specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
@@ 3 Dataset
We investigate the climate campaigns case study for this work. We work on the corpus of $21372$ ads Islam and Goldwasser (2024a) originally released by Islam et al. (2023b). This dataset includes climate-related English ads on Facebook from the US, spanning from January $2021$ to January $2022$. Each ad includes the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, and distribution of impressions broken down by gender (male, female, unknown), age (seven groups), and location down to the state level in the USA. Additional details about the dataset can be found in the original publications.
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). However, Facebook Ad Library API doesn’t provide explicit targeting information beyond ad impressions (views). We focus only on ads viewed exclusively by either males or females; ads viewed by both genders were excluded. Therefore, if an ad is viewed solely by males or solely by females, we consider it as a targeted ad. The same logic is applied to age groups. We do not include any ads that overlap across age categories such as young adults, early working-age individuals, late working-age individuals, and senior citizens. By following this methodology, we narrow down the dataset to $227$ ads. Among them, $59$ ads targeting only females and $47$ ads targeting only males. However, we find $25$ ads target only young adults, $82$ ads target only the early working age group, $8$ ads target only the late working age group, and $6$ ads target only senior citizens.
@@ 4.1 Results
o1-preview age 85.95 71.00
Table 2: Baseline comparisons. Demo. : Demographics.
@@ 4.2 Statistical Tests
Senior $4.66\times 10^{-1}$ Fail to Reject H0
@@ 5 Thematic Insights of Explanations
As LLMs provide explanations to provide reasoning behind their prediction, we use those explanations for understanding thematic insights to answer RQ2. In this analysis, we only include the correct predictions and their explanations.
@@ 5.3 Baselines
We add Latent Dirichlet Allocation (LDA) Blei et al. (2003) and BERTopic Grootendorst (2022) as unsupervised techniques for generating topics from LLMs-generated explanations. BERTopic [^8] https://maartengr.github.io/BERTopic/#common is a hierarchical topic modeling and an off-the-shelf neural baseline for clustering texts. BERTopic generates two topics from the prediction explanation of gender and age (Table 7). We run LDA with four topics and show the top $10$ words in Table 7.
Dimension Method Topics / Top Words
@@ 6 Fairness and Bias Analysis
Ad Text Prediction Explanations by LLMs
Table 8: Misclassified Senior Instances.
@@ 6.1 Fairness Analysis on Gender Prediction
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by $0.10$, indicating that model may be more effective at correctly identifying females than males
@@
Table 11: Equal Opportunity (True Positive Rate) and Predictive Equality (False Positive Rate) for Age Groups.
@@ 6.2 Fairness Analysis on Age Prediction
To understand the observed bias in the misclassified senior age group, we conduct an analysis of the misclassified instances (Table 8). The goal is to identify patterns and underlying reasons why the model predicts seniors as belonging to the young and late working age group. The misclassification of the Senior age group as Young or Late Working can be attributed to several factors:
@@ 7 Conclusion
Our work advances the discourse on climate microtargeting by demonstrating the utility of LLMs in accurately detecting and explaining targeted messaging strategies on social media. We show our analysis on climate campaigns dataset, but our approach can easily be adapted in any dataset. It is designed to be scalable without any modifications. We provide a broader theme and various aspects under each theme based on the explanations from LLMs, which we hope will be an important contribution to the CSS community. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates, particularly in underrepresented groups, highlight the need for more inclusive and equitable targeting methods. Ultimately, this study lays the groundwork for future investigations into the role of AI in enhancing the efficacy, transparency, and accountability of digital climate advocacy.
@@ 8 Limitations
The original dataset does not specify intersectional targeting, such as young females only. Our paper acknowledges that our approach infers targeting from exclusive impression data (view), since the Facebook Ad Library API does not disclose actual targeting parameters.
We have small sample size due to restrictive filtering criteria, might result in underrepresented groups, particularly late working and senior demographics.

### v3

@@ Abstract
Climate change communication on social media increasingly employs microtargeting strategies to effectively reach and influence specific demographic groups. This study presents a post-hoc analysis of microtargeting practices within climate campaigns by leveraging large language models (LLMs) to examine Meta (previously known as Facebook) advertisements. Our analysis focuses on two key aspects: demographic targeting and fairness. We evaluate the ability of LLMs to accurately predict the intended demographic targets, such as gender and age group. Furthermore, we instruct the LLMs to generate explanations for their classifications, providing transparent reasoning behind each decision. These explanations reveal the specific thematic elements used to engage different demographic segments, highlighting distinct strategies tailored to various audiences. Our findings show that young adults are primarily targeted through messages emphasizing activism and environmental consciousness, while women are engaged through themes related to caregiving roles and social advocacy. Additionally, we conduct a comprehensive fairness analysis to uncover biases in model predictions. We assess disparities in accuracy and error rates across demographic groups using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality. Our findings indicate that while LLMs perform well overall, certain biases exist, particularly in the classification of male audiences. The analysis of thematic explanations uncovers recurring patterns in messaging strategies tailored to various demographic groups, while the fairness analysis underscores the need for more inclusive targeting methods. This study provides a valuable framework for future research aimed at enhancing transparency, accountability, and inclusivity in social media-driven climate campaigns.
@@ 2 Related Work
Microtargeting on social media platforms such as Meta has been widely studied in the context of political campaigns Islam et al. (2023a); Capozzi et al. (2021); Serrano et al. (2020); Silva et al. (2020); Capozzi et al. (2020) and public health messaging Islam and Goldwasser (2022); Silva and Benevenuto (2021); Mejova and Kalimeri (2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups Islam et al. (2023b). However, the effectiveness and ethical implications of such targeted messaging remain under-explored, particularly in terms of how different demographic groups are engaged and whether biases are present in the targeting strategies.
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans Chowdhery et al. (2023); Kojima et al. (2022); Le Scao et al. (2022); Brown et al. (2020); Gilardi et al. (2023); De Paoli (2023); Dai et al. (2023); Chiang and Lee (2023); Ziems et al. (2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into Thematic Analysis (TA). Researchers have proposed various frameworks, including an LLMs-in-the-loop approach Islam and Goldwasser (2024b, a); Dai et al. (2023), integrating GPT-3 with expert-designed codebooks Xiao et al. (2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes Gao et al. (2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer Wei et al. (2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs predictions which have been highlighted in recent literatures Anthis et al. (2024); Dai et al. (2024); Kumar et al. (2023); Li et al. (2023b); Wu et al. (2023); Bender et al. (2021), emphasizing the need for rigorous evaluation of biases and disparities in model performance across different groups. In this paper, we leverage LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Meta. Besides, we develop a new set of themes and aspects based on those explanations specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
@@ 3 Dataset
We investigate the climate campaigns case study for this work. We work on the corpus of $21372$ ads Islam and Goldwasser (2024a) originally released by Islam et al. (2023b). This dataset includes climate-related English ads on Meta from the US, spanning from January $2021$ to January $2022$. Each ad includes the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, and distribution of impressions broken down by gender (male, female, unknown), age (seven groups), and location down to the state level in the USA. Additional details about the dataset can be found in the original publications.
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). However, Meta Ad Library API doesn’t provide explicit targeting information beyond ad impressions (views). Note that this paper infers targeting from exclusive impression distributions, which may differ from advertiser intent. In this work, we focus only on ads viewed exclusively by either males or females; ads viewed by both genders were excluded. Therefore, if an ad is viewed solely by males or solely by females, we consider it as a targeted ad. The same logic is applied to age groups. We do not include any ads that overlap across age categories such as young adults, early working-age individuals, late working-age individuals, and senior citizens. By following this methodology, we narrow down the dataset to $227$ ads. Among them, $59$ ads targeting only females and $47$ ads targeting only males. However, we find $25$ ads target only young adults, $82$ ads target only the early working age group, $8$ ads target only the late working age group, and $6$ ads target only senior citizens.
@@ 4.1 Results
age 85.95 71.00
Table 2: Baseline comparisons. Demo.: Demographics, Acc.: Accuracy.
@@ 4.2 Statistical Tests
Senior $4.66\times 10^{-1}$ Fail Reject H0
@@ 5 Thematic Insights of Explanations
As LLMs provide explanations to provide the reasoning behind their prediction, we use those explanations to understand thematic insights to answer RQ2. In this analysis, we only include the correct predictions and their explanations. Using only correct predictions is not arbitrary—it is supported by methodological norms (Vaismoradi et al., 2016, 2013; Braun and Clarke, 2012, 2006) to ensure pattern coherence.
@@ 5.3 Baselines
We add Latent Dirichlet Allocation (LDA) Blei et al. (2003) and BERTopic Grootendorst (2022) as unsupervised techniques for generating topics from LLMs-generated explanations. BERTopic [^8] https://maartengr.github.io/BERTopic/#common is a hierarchical topic modeling and an off-the-shelf neural baseline for clustering texts. BERTopic generates two topics from the prediction explanation of gender and age (Table 13 in App. A.5). We run LDA with four topics and show the top $10$ words in App. A.5 in Table 13. We perform a human evaluation by asking annotators to assess whether the generated themes and topics match with ad’s explanation. Table 12 in App. A.5 shows the human evaluation results on LLM (GPT-4o OpenAI (2024)) assignment, LDA, and BERTopic.
@@ 6.1 Fairness Analysis on Gender Prediction
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by $0.10$, indicating that model may be more effective at correctly identifying females than males.
@@
Table 9: Equal Opportunity (TPR) and Predictive Equality (FPR) for Age Groups.
@@ 6.2 Fairness Analysis on Age Prediction
To understand the observed bias in the misclassified senior age group, we conduct an analysis of the misclassified instances (Table 11 in App. A.2.1). The goal is to identify patterns and underlying reasons why the model predicts seniors as belonging to the young and late working age group. The misclassification of the Senior age group as Young or Late Working can be attributed to several factors, such as: lack of age-specific cues, reliance on stereotypical associations, feature representation limitations etc. We detail in App. A.2.1.
@@ 6.3 Statistical Significance Tests
6.4 Sensitivity Analysis
Our study is descriptive and does not claim causal effects. We have added sensitivity analysis (Fig. 7 in App. A.6) by rerunning the fairness test after stratifying by ad-spend and by topic cluster to confirm the reported biases persist under two plausible confounders.
Fig 7(a) in App. A.6 shows that bias persists after conditioning on ad spend. Low-budget ads: $\Delta\approx 0.82$ ($95\%$ CI 0.680.680.680.680.960.960.960.96), High-budget ads: $\Delta\approx 0.76$ ($95\%$ CI 0.550.550.550.550.930.930.930.93). Overlapping CIs indicate similar magnitude and ad-budget is not driving the disparity. Topic clusters tell the same story (Fig 7(b) in App. A.6). Four of five clusters show sizable positive gaps (0.300.300.300.301.001.001.001.00). The one cluster with only female-authored ads (topic 2) cannot yield a gap, but its omission does not overturn the pattern. Accuracy gaps are modest and hover around zero. Where CIs cross zero, classification accuracy is comparable between genders; the selection-rate bias remains the dominant concern.
[…]
6.5 Robustness Analysis
To provide a more in-depth evaluation of the quality and robustness of the data, we plot learning curve containing macro-F1 and demographic-parity against sample size (20202020227227227227 ads). Learning-curve plot in Fig. 8 in App. A.7 shows that macro-F1 and fairness gaps plateau well before 227 examples—evidence that the analysis is not variance-limited.
Both curves (Fig. 8(a) and Fig. 8(b) App. A.7) plateau by $\approx 60$ examples, confirming that our results are robust to subsampling and that additional data in the current distribution yield diminishing returns. Both performance and bias saturate after $\sim 1/4$ of the data, indicating the analysis is not variance-limited.
Figures 8(c) and 8(d) in App. A.7 show the macro-F1 and age-parity gap at 20-ad increments. Macro-F1 rises to $0.70$ by $n\approx 80$ and then plateaus; the $95\%$ confidence band narrows to ±$0.8$ pp. The parity gap likewise converges to $\approx 0.57$ with error $<0.02$. Thus, our age-bias findings are not driven by sample variance—adding more data in the current distribution neither boosts accuracy nor mitigates disparity.
@@ 7 Conclusion
Our work advances the discourse on climate microtargeting by demonstrating the utility of LLMs in accurately detecting and explaining targeted messaging strategies on social media. We show our analysis on climate campaigns dataset, but our approach can easily be adapted in any dataset. It is designed to be scalable without any modifications. We provide a broader theme and various aspects under each theme based on the explanations from LLMs, which we hope will be an important contribution to the CSS community. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates, highlight the need for more inclusive and equitable targeting methods.
@@ 8 Limitations
The original dataset does not specify intersectional targeting, such as young females only. Our paper acknowledges that our approach infers targeting from exclusive impression data (view), since the Meta Ad Library API does not disclose actual targeting parameters. Meta’s own transparency studies treat ads whose impressions are $\geq 95\%$ in one group as “effectively exclusive". Such ads deliver the same real-world exposure as explicit targeting.
We have small sample size due to restrictive filtering criteria which might result in underrepresented groups, particularly late working and senior demographics. The rationale for using this specific dataset is that it is publicly available and, to the best of our knowledge, it is the only publicly available dataset focused on targeted climate campaigns.
@@ A.2 Demographic Misclassifications
A.2.1 Misclassification - Senior Instances
Table 11 shows the details of misclassified senior instances. The misclassification of the Senior age group as Young or Late Working can be attributed to several factors:
Ad Text Prediction Explanations by LLMs
Table 11: Misclassified Senior Instances.
@@ A.4 Thematic Analysis of Age Explanations
A.5 Unsupervised Baselines
Generated topics from prediction explanation of gender and age using LDA and BERTopic is shown in Table 13. Table 12 shows the human evaluation results on GPT-4o assignment, LDA, and BERTopic. Prompt template for LLM assignment is shown in Fig. 6.
Model Demo. Acc. (%)
LDA gender 61.32
age 33.06
BERTopic gender 56.60
age 45.45
GPT-4o gender 88.67
age 68.56
Table 12: Assignment Comparison w.r.t. Human Judgments. Demo.: Demographics, Acc.: Accuracy.
Dimension Method Topics / Top Words
Figure 6: Prompt template for assigning theme using LLM. Inputs are shown in blue, and outputs are shown in red.
A.6 Sensitivity Analysis on Fairness
We explicitly frame our work as a post-hoc audit and the manuscript avoids causal verbs such as “impact” or “cause” when reporting accuracy or fairness metrics. Fig. 7 shows the sensitivity analysis by rerunning the fairness test after stratifying by ad-spend and by topic cluster.
(a) Fairness by spend tier (Gender).
(b) Fairness by topic cluster (Gender).
(c) Fairness by spend tier (Age)
(d) Fairness by topic cluster (Age).
Figure 7: Sensitivity Analysis: Rerunning the fairness test after stratifying by ad-spend and by topic cluster.
A.7 Learning-Curve Robustness Analysis
Fig. 8 shows the learning curve having macro average F1 score vs. sample size and fairness gaps vs. sample size.
(a) F1 vs sample size (Gender).
(b) Parity gap vs sample size (Gender).
(c) F1 vs sample size (Age).
(d) Parity gap vs sample size (Age).
Figure 8: Learning-curve robustness analysis.